The mean lifetime of a sample of 144 fluorescent light bulbs produced by a company is computed to be 170 hours with a standard deviation of 120 hours. Test the company's claim that its bulbs' mean life time is 200 hours at 0.05 significance level. H0 : = 1600 hours H1 : 1600 hours

1
Expert's answer
2021-09-20T16:42:06-0400

"\bar{x} = 170 \\

s = 120 \\

n=144 \\

H_0: \mu = 200 \\

H_1: \mu ≠200"

Test-statistic:

"t=\frac{\bar{x} -\mu}{s/ \sqrt{n}} \\

t = \frac{170-200}{120 / \sqrt{144}} =\frac{-30}{10} = -3 \\

α= 0.05 \\

df=n-1 = 144-1 = 143"

Two-tailed test

From table:

"t_{\u03b1\/2,df} = t_{0.05\/2, 143} = +\/-1.9767"

The decision rule is: Reject H0 if t ≤ -1.9767 or if t ≥ 1.9767.

Since, "t=-3< t_{\u03b1\/2,df}"

We reject H0 at 0.05 level of significance.

Conclusion: There is NOT sufficient evidence to support the claim that "\mu = 200."